June is 15 years older than Chuck is now. 5 years from now June will be twice as old as Chuck. Find their ages now.
I used a number line to get the answer. 25 June. 10 Chuck. I can't figure out the equation.

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39616)
Answer by ikleyn(52776)

.
June is 15 years older than Chuck is now. 5 years from now June will be twice as old as Chuck. Find their ages now.
I used a number line to get the answer. 25 June. 10 Chuck. I can't figure out the equation.
~~~~~~~~~~~~~~~~~

Now:               J = C + 15.     (1)   ("June is 15 years older than Chuck is now. ")

5 years from now:  J+5 = 2*(C+5)   (2)   ("5 years from now June will be twice as old as Chuck.")


Substitute (1) into (2). You will get

(C+15) + 5 = 2(C+5),

C + 20 = 2C + 10,

20 - 10 = 2C - C  --->  C = 10.

Answer. Chuck is 10 years old now. June is 10+15 = 25 years now.


There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".